OpenStudy (calculusxy):

1. How far will an object move in one second if its average speed is 5 m/s? A. 5 meters 2. How far will a freely falling object have fallen from a position of rest when its instantaneous speed is 10 m/s? A. 75 meters 3. An object dropped from rest and falls freely. After 6 seconds, calculate its instantaneous speed, average speed, and distance fallen. A. Instantaneous speed: 58.8 m/s Average Speed : 53.9 m/s Distance : 313.6 meters 4. If a freely falling rock were equipped with an odometer, would the readings for the distance fallen each second stay the same...

OpenStudy (calculusxy):

...increase with tie, or decrease with time? A. The instantaneous speed would increase so the distance would increase as well.

OpenStudy (calculusxy):

@jim_thompson5910

OpenStudy (calculusxy):

Can you make sure the answers?

OpenStudy (ybarrap):

A. Is correct because $$ 5 \frac{m}{\cancel{s}}\times 1~\cancel{s}=5~m $$ Notice how the seconds cancel each other out and you are left with just meters ,\(m\)?

OpenStudy (calculusxy):

Yeah

OpenStudy (ybarrap):

For 2 you need to use the following equation $$ v^2 = u^2 + 2as $$ where u is the initial velocity, a is the acceleration of gravity, s is the distance traveled and v is the final velocity: https://en.wikipedia.org/wiki/Equations_of_motion#Kinematic_equations_for_one_particle So $$ 10^2=0^2+2s\times 9.81\\ \implies s=\frac{100}{2\times 9.81}=5~m $$ Does this make sense?

OpenStudy (calculusxy):

Answer would be 5M?

OpenStudy (ybarrap):

Yes

OpenStudy (calculusxy):

how would it be 5m if it is travelling 10m/s

OpenStudy (ybarrap):

 (ignore the t=1/2 seconds above) It takes $$ v/a=10/9.81=1.01\text{ second} $$ to go from 0 to 10 m/s In that time period, it is accelerating. If it were going a constant speed of 10 m/s then is 1 second it would have traveled 10 m, but it didn't START at 10 m/s, it started at 0 m/s. Right?